OpenStudy (anonymous):

Solve by elimination 3x - 4y = 20 5x + 8y = 48 I got (8,1) is that correct

OpenStudy (bahrom7893):

let's check: 3*8-4*1 = 24 - 4 = 20 (correct) 5*8+8*1=40+8=48 (correct)

OpenStudy (bahrom7893):

You are right

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

Im going to see if more of mines are right so im going to post them then type your name so u can get medal that cool

OpenStudy (anonymous):

Thx

OpenStudy (bahrom7893):

np